# Scaling and Load-Balancing Equi-Joins

### v2

@@ 3.1.2 Shared-Nothing Equi-Joins
The algorithm in [45] is similar to that in [65] discussed above. It allows for grouping the records into evenly distributed groups that are executed on individual executors. However, it suffers from the same hefty overhead.
@@
Map-Reduce-Merge [93] extends MapReduce with a merge function to facilitate expressing the join operation. Map-Join-Reduce [49] adds a join phase between the map and the reduce phases. These extensions cannot leverage the MapReduce built-in fault tolerance, and are implemented using complicated custom logic.
@@
The Partial Redistribution & Partial Duplication (PRPD) algorithm in [92] is a hybrid between hash-based [52, 38] (as opposed to Sort-Merge [42, 10, 14]) and duplication-based join algorithms [39]. The PRPD algorithm collects $\kappa_{\mathcal{R}}$, and $\kappa_{\mathcal{S}}$, the hot (a.k.a. high-frequency) keys in $\mathcal{R}$ and $\mathcal{S}$, respectively. For the correctness of the algorithm, a key that is hot in both $\mathcal{R}$ and $\mathcal{S}$ is only assigned to either $\kappa_{\mathcal{R}}$ or to $\kappa_{\mathcal{S}}$. PRPD splits $\mathcal{S}$ into (a) $\mathcal{S}_{high}$ with keys in $\kappa_{\mathcal{S}}$, whose records are not distributed to the executors, (b) $\mathcal{S}_{dup}$ with keys in $\kappa_{\mathcal{R}}$, whose records are broadcasted to all the executors containing $\mathcal{R}_{high}$ records, and (c) $\mathcal{S}_{hash}$ the remaining records that are distributed among executors using hashing. $\mathcal{R}$ is split similarly into $\mathcal{R}_{high}$, $\mathcal{R}_{dup}$, and $\mathcal{R}_{hash}$. $\mathcal{S}_{hash}$ is joined with $\mathcal{R}_{hash}$ on the executors they were hashed into, and $\mathcal{S}_{high}$ ($\mathcal{R}_{high}$) is joined with $\mathcal{R}_{dup}$ ($\mathcal{S}_{dup}$) on the executors containing $\mathcal{S}_{high}$ ($\mathcal{R}_{high}$). The union of the three joins comprise the final join results [^5] The unioning of datasets in a distributed system is typically executed by unioning the sets of files containing them, where each file is specified by both the executor it resides on, and the path to the file on that matchine..
@@
Track-Join [68] aims at minimizing the network load using a framework similar to PRPD. For any given key, Track-Join migrates its records from one relation to a few executors [69], and selectively broadcasts, i.e., multicasts, the records from the other relation to these executors. This record migration process is expensive. Moreover, identifying the executors on which the join takes place is done in a preprocessing track phase, which is a separate distributed join that is sensitive to skew in key popularity.
@@
The PRPD algorithm is developed further into the SkewJoin algorithm in [21]. Three flavors of SkewJoin are proposed. The Broadcast-SkewJoin (BSJ) is the same as PRPD. On the other end of the spectrum, Full-SkewJoin (FSJ) distributes $\mathcal{S}_{high}$ and $\mathcal{R}_{high}$ in a round-robin fashion on the executors using a specialized partitioner. This offers better load balancing at the cost of higher distribution overhead. The Hybrid-SkewJoin (HSJ) is a middle ground that retains the original data partitioning on the non-skewed side.
@@
The PRPD algorithm is extended by [26] to use distributed lookup servers [^6] Utilization lookup servers has been proposed before by the same research group [25].. However, the algorithm in [26] splits $\mathcal{S}$ into $\mathcal{S}_{hash}$ and $\mathcal{S}_{high}$. $\mathcal{S}_{hash}$, along with the distinct keys of $\mathcal{S}_{high}$, are hashed into the executors. All $\mathcal{R}$ is hashed into the executors. On each executor, $e_{i}$, $\mathcal{R}_{i}$, the partition of $\mathcal{R}$ on $e_{i}$, is joined with $\mathcal{S}_{hash_{i}}$, and with $\mathcal{S}_{high_{i}}$. The results of joining $\mathcal{R}_{i}$ with $\mathcal{S}_{hash_{i}}$ are reported as part of the final results. The results of joining $\mathcal{R}_{i}$ with $\mathcal{S}_{high_{i}}$ are augmented with the right $\mathcal{S}$ records using the distributed lookup servers.
@@
Flow-Join [71] is similar to PRPD, but it detects the hot keys while performing the cross-network join. Flow-Join utilizes (a) Remote Direct Memory Access (RDMA) [55] over high-speed network [72, 17, 74] and (b) work-stealing across cores and Non-Uniform Memory Access (NUMA) [72] regions for local processing on rack-scale systems.
3.1.3 Comments on the General Distributed Equi-Joins
@@
All the above algorithms that mix Broadcast-Joins and Shuffle-Joins [92, 68, 69, 21, 26, 71] do not extend smoothly to the outer-join variants. Mixing these two join techniques in the PRPD fashion results in having dangling tuples (tuples that do not join on the other side) of the same key distributed across the network. Deduplicating and counting dangling tuples across the network is a non-trivial operation, and was only discussed in [21, 71].
The single-executor-per-key Shuffle-Join (Hash or Sort-Merge) algorithms that assume the Cartesian product of the records of any given key is computed using one executor fail to handle highly skewed data, where one key can bottleneck an entire executor. Consequently, all the other executors remain under-utilized while waiting for the bottleneck executor to finish [84].
On the other hand, the multi-executor-per-key Shuffle-Join algorithms that distributes the work of one hot key on multiple executors in [2, 65, 15, 56] have two major drawbacks. First, computing a decent assignment of the executors to the grids of the hot keys is fairly costly. The most optimized work in [56] balances the load between the executors in a near-optimal way as described above and under the assumptions stated above. Estimating the variance is $\Theta(n\log(n))$ for increasing a grid dimension by one executor, where $n$ is the number of executors. This is done for all hot keys to find the best reduction in variance, which is $\Theta(\min(|\kappa_{\mathcal{R}}|_{max}$, $|\kappa_{\mathcal{S}}|_{max}))$, where $|\kappa_{\mathcal{R}}|_{max}$ and $|\kappa_{\mathcal{S}}|_{max}$ are the maximum numbers of hot keys collected from $\mathcal{R}$ and $\mathcal{S}$, respectively. Finally, there are $\Omega(n)$ iterations executed to stabilize the expected run time. Hence, the work done by the driver before assigning hot keys to executors is $\Omega(n^{2}\log(n))\times\min(|\kappa_{\mathcal{R}}|_{max}$, $|\kappa_{\mathcal{S}}|_{max})$.
@@
The strongest drawback of all the algorithms that use Broadcast-Join is low scalability when there exist keys that are hot in both joined relations. In such cases, the broadcasted keys are hot, and their records may not fit in memory. Even if the records of the hot keys may fit in memory, the entire broadcasted relation may not fit in memory. The discussion assumes the basic Broadcast-Join implementation. The Broadcast-Join algorithm can be made more scalable by pre-splitting $\mathcal{S}$ into multiple sub-relations, $\mathcal{S}_{1},\dots,\mathcal{S}_{i},\dots\mathcal{S}_{m}$, for some $m$, such that each sub-relation, $\mathcal{S}_{i}$ fits in memory, running a Broadcast-Join between $\mathcal{R}$ and each sub-relation $\mathcal{S}_{i}$, and unioning the results of all the joins. However, the standard implementations of Broadcast-Join in major DBMSs are not known to provide this partition and broadcast [56] optimization by default.
3.2 Small-Large Outer-Joins
@@
For a left-outer-join, the Duplication and Efficient Redistribution (DER) algorithm [91] broadcasts the small relation, $\mathcal{S}$, to all $n$ executors. On each executor, $e_{i}$, $\mathcal{S}$ is inner joined with $\mathcal{R}_{i}$. The joined records are output, and the ids of the unjoined records of $\mathcal{S}$ are distributed to the $n$ executors based on their hash. If an executor receives $n$ copies of a record id, then this $\mathcal{S}$ id is unjoinable with $\mathcal{R}$ (since it was unjoined on the $n$ executors). The unjoinable ids are inner Hash-Joined with $\mathcal{S}$ and $null$-padded. The left-outer-join result is the union of the two inner join results.
The Duplication and Direct Redistribution (DDR) algorithm [27] is similar to DER, but for the records failing the first inner join, each executor hashes the entire unjoined record, instead of only its id, over the network. Because entire records, instead of their ids, are hash-distributed, (a) the first join can be executed as an out-of-the-box left-outer-join, and (b) the second inner join is not needed. Extending DER and DDR to full-outer-joins is achieved by changing the first join to output the unjoined records on $\mathcal{R}_{i}$.
4 The Tree-Join Algorithm
@@
4.1 The Basic Tree-Join Algorithm
@@
4:The secod set of lists are the chunks of $\mathcal{L}_{2}$.
@@
Figure 3: A schematic example for joining a specific key, $k_{1}$, using the basic Tree-Join algorithm. (a) The $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The $k_{1}$ joined lists in the initial joined index formed using groupByKey. Each of the joined lists are chunked into two sub-lists (to simplify the figure). (c) Each sub-list from $\mathcal{R}$ are produced with all the sub-lists from $\mathcal{S}$. These pairs of sublists can be processed by different executors in the next iteration. (b) and (c) together show the top two levels of the tree formed by the basic Tree-Join algorithm.
This tree of the execution iterations is illustrated in Fig. 3 for an example key, $k_{1}$. Fig. 3(a) shows the $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. Fig. 3(b) shows the basic Tree-Join algorithm forming the joined lists of $k_{1}$ in the initial joined index, $joined\_index_{0}$, which is the top level of the tree. The joined lists are chunked into sub-lists. To simplify Fig. 3, each list is chunked into two sub-lists. That is, $\mathcal{L}_{\mathcal{R}1}$ is chunked into $\mathcal{L}_{\mathcal{R}11}$ and $\mathcal{L}_{\mathcal{R}12}$, and $\mathcal{L}_{\mathcal{S}1}$ is chunked into $\mathcal{L}_{\mathcal{S}11}$ and $\mathcal{L}_{\mathcal{S}12}$. Fig. 3(c) shows the basic Tree-Join algorithm forming the second joined index, $joined\_index_{1}$, which is the next level of the tree. This is done by outputting each sub-list from $\mathcal{R}$, $\mathcal{L}_{\mathcal{R}11}$ and $\mathcal{L}_{\mathcal{R}12}$, with all the sub-lists from $\mathcal{S}$, $\mathcal{L}_{\mathcal{S}11}$ and $\mathcal{L}_{\mathcal{S}12}$. Notice these joined sub-lists have a repeated key in $joined\_index_{1}$ Hence, these keys can can be distributed on different data partitions and hence processed by different executors for load balancing. To simplify Fig. 3, only the first two levels of the execution iteration tree are shown in Fig. 3(b) and Fig. 3(c), respectively.
Figure 4: An example of inner-joining two relations using treeJoinBasic. (a) The input relations. (b) The initial index built from the input relations in (a). The initial index is used to produce (c) the first partial results, and (d) the first joined index. The first joined index is used to produce (e) the second partial results, and (f) the second joined index. The second joined index is used to produce (g) the third partial results. The union of the partial results in (c), (e), and (g) constitutes the inner-join results.
@@
4.2 Chunking Lists of Hot Keys
As explained in § 4.1, The treeJoinBasic algorithm executes multiple iterations. During each iteration, for each hot key, its joined lists are chunked by the splitter, the executor assigned the key and its joined lists, into sub-lists to be processed by the executors of the subsequent iteration. The subsequent executors can (a) output the pairs of records in the joined sub-lists, or (b) act as further splitters of the sub-lists if they are still too long. The chunking of a joined sub-lists continues until they are short enough.
Computing the optimal number of sub-lists for a list of records is crucial for load balancing between the executors and across iterations. Chunking a pair of joined lists into numerous pairs of sub-lists adds overhead to the splitter. Conversely, chunking the pair of joined lists into few pairs of sub-lists adds overhead to the executors that handle these pairs in the subsequent iteration.

### v3

@@ 3.1.2. Shared-Nothing Equi-Joins
The algorithm in (Hassan and Bamha, 2009) is similar to that in (Okcan and Riedewald, 2011) discussed above. It allows for grouping the records into evenly distributed groups that are executed on individual executors. However, it suffers from the same hefty overhead.
@@
Map-Reduce-Merge (Yang et al., 2007) extends MapReduce with a merge function to facilitate expressing the join operation. Map-Join-Reduce (Jiang et al., 2010) adds a join phase between the map and the reduce phases. These extensions cannot leverage the MapReduce built-in fault tolerance, and are implemented using complicated custom logic.
@@
The Partial Redistribution & Partial Duplication (PRPD) algorithm in (Xu et al., 2008) is a hybrid between hash-based (Kitsuregawa et al., 1983; DeWitt et al., 1990) (as opposed to Sort-Merge (Graefe, 1994; Balkesen et al., 2013b; Barthels et al., 2017)) and duplication-based join algorithms (DeWitt et al., 1992). The PRPD algorithm collects $\kappa_{\mathcal{R}}$, and $\kappa_{\mathcal{S}}$, the hot (a.k.a. high-frequency) keys in $\mathcal{R}$ and $\mathcal{S}$, respectively. For the correctness of the algorithm, a key that is hot in both $\mathcal{R}$ and $\mathcal{S}$ is only assigned to either $\kappa_{\mathcal{R}}$ or to $\kappa_{\mathcal{S}}$. PRPD splits $\mathcal{S}$ into (a) $\mathcal{S}_{high}$ with keys in $\kappa_{\mathcal{S}}$, whose records are not distributed to the executors, (b) $\mathcal{S}_{dup}$ with keys in $\kappa_{\mathcal{R}}$, whose records are broadcasted to all the executors containing $\mathcal{R}_{high}$ records, and (c) $\mathcal{S}_{hash}$ the remaining records that are distributed among executors using hashing. $\mathcal{R}$ is split similarly into $\mathcal{R}_{high}$, $\mathcal{R}_{dup}$, and $\mathcal{R}_{hash}$. $\mathcal{S}_{hash}$ is joined with $\mathcal{R}_{hash}$ on the executors they were hashed into, and $\mathcal{S}_{high}$ ($\mathcal{R}_{high}$) is joined with $\mathcal{R}_{dup}$ ($\mathcal{S}_{dup}$) on the executors containing $\mathcal{S}_{high}$ ($\mathcal{R}_{high}$). The union of the three joins comprise the final join results [^5] The unioning of datasets in a distributed system is typically executed by unioning the sets of files containing them, where each file is specified by both the executor it resides on, and the path to the file on that matchine..
@@
Track-Join (Polychroniou et al., 2014) aims at minimizing the network load using a framework similar to PRPD. For any given key, Track-Join migrates its records from one relation to a few executors (Polychroniou et al., 2018), and selectively broadcasts, i.e., multicasts, the records from the other relation to these executors. This record migration process is expensive. Moreover, identifying the executors on which the join takes place is done in a preprocessing track phase, which is a separate distributed join that is sensitive to skew in key popularity.
@@
The PRPD algorithm is developed further into the SkewJoin algorithm in (Bruno et al., 2014). Three flavors of SkewJoin are proposed. The Broadcast-SkewJoin (BSJ) is the same as PRPD. On the other end of the spectrum, Full-SkewJoin (FSJ) distributes $\mathcal{S}_{high}$ and $\mathcal{R}_{high}$ in a round-robin fashion on the executors using a specialized partitioner. This offers better load balancing at the cost of higher distribution overhead. The Hybrid-SkewJoin (HSJ) is a middle ground that retains the original data partitioning on the non-skewed side.
@@
The PRPD algorithm is extended by (Cheng et al., 2014) to use distributed lookup servers [^6] Utilization of lookup servers has been proposed before by the same research group (Cheng et al., 2013).. However, the algorithm in (Cheng et al., 2014) splits $\mathcal{S}$ into $\mathcal{S}_{hash}$ and $\mathcal{S}_{high}$. $\mathcal{S}_{hash}$, along with the distinct keys of $\mathcal{S}_{high}$, are hashed into the executors. All $\mathcal{R}$ is hashed into the executors. On each executor, $e_{i}$, $\mathcal{R}_{i}$, the partition of $\mathcal{R}$ on $e_{i}$, is joined with $\mathcal{S}_{hash_{i}}$, and with $\mathcal{S}_{high_{i}}$. The results of joining $\mathcal{R}_{i}$ with $\mathcal{S}_{hash_{i}}$ are reported as part of the final results. The results of joining $\mathcal{R}_{i}$ with $\mathcal{S}_{high_{i}}$ are augmented with the right $\mathcal{S}$ records using the distributed lookup servers.
@@
Flow-Join (Rödiger et al., 2016) is similar to PRPD, but it detects the hot keys while performing the cross-network join. Flow-Join utilizes (a) Remote Direct Memory Access (RDMA) (Li et al., 2016) over high-speed network (Rödiger et al., 2015; Binnig et al., 2016; Salama et al., 2017) and (b) work-stealing across cores and Non-Uniform Memory Access (NUMA) (Rödiger et al., 2015) regions for local processing on rack-scale systems.
3.1.3. Comments on the General Distributed Equi-Joins
@@
All the above algorithms that mix Broadcast-Joins and Shuffle-Joins (Xu et al., 2008; Polychroniou et al., 2014, 2018; Bruno et al., 2014; Cheng et al., 2014; Rödiger et al., 2016) do not extend smoothly to the outer-join variants. Mixing these two join techniques in the PRPD fashion results in having dangling tuples (tuples that do not join on the other side) of the same key distributed across the network. Deduplicating and counting dangling tuples across the network is a non-trivial operation, and was only discussed in (Bruno et al., 2014; Rödiger et al., 2016).
The single-executor-per-key Shuffle-Join (Hash or Sort-Merge) algorithms that assume the Cartesian product of the records of any given key is computed using one executor fail to handle highly skewed data, where one key can bottleneck an entire executor. Consequently, all the other executors remain under-utilized while waiting for the bottleneck executor to finish (Suri and Vassilvitskii, 2011).
On the other hand, the multi-executor-per-key Shuffle-Join algorithms that distributes the work of one hot key on multiple executors in (Afrati and Ullman, 2010; Okcan and Riedewald, 2011; Beame et al., 2014; Li et al., 2018) have two major drawbacks. First, computing a decent assignment of the executors to the grids of the hot keys is fairly costly. The most optimized work in (Li et al., 2018) balances the load between the executors in a near-optimal way as described above and under the assumptions stated above. Estimating the variance is $\Theta(n\log(n))$ for increasing a grid dimension by one executor, where $n$ is the number of executors. This is done for all hot keys to find the best reduction in variance, which is $\Theta(\min(|\kappa_{\mathcal{R}}|_{max}$, $|\kappa_{\mathcal{S}}|_{max}))$, where $|\kappa_{\mathcal{R}}|_{max}$ and $|\kappa_{\mathcal{S}}|_{max}$ are the maximum numbers of hot keys collected from $\mathcal{R}$ and $\mathcal{S}$, respectively. Finally, there are $\Omega(n)$ iterations executed to stabilize the expected run time. Hence, the work done by the driver before assigning hot keys to executors is $\Omega(n^{2}\log(n))\times\min(|\kappa_{\mathcal{R}}|_{max}$, $|\kappa_{\mathcal{S}}|_{max})$.
@@
The strongest drawback of all the algorithms that use Broadcast-Join is low scalability when there exist keys that are hot in both joined relations. In such cases, the broadcasted keys are hot, and their records may not fit in memory. Even if the records of the hot keys may fit in memory, the entire broadcasted relation may not fit in memory. The discussion assumes the basic Broadcast-Join implementation. The Broadcast-Join algorithm can be made more scalable by pre-splitting $\mathcal{S}$ into multiple sub-relations, $\mathcal{S}_{1},\dots,\mathcal{S}_{i},\dots\mathcal{S}_{m}$, for some $m$, such that each sub-relation, $\mathcal{S}_{i}$ fits in memory, running a Broadcast-Join between $\mathcal{R}$ and each sub-relation $\mathcal{S}_{i}$, and unioning the results of all the joins. However, the standard implementations of Broadcast-Join in major DBMSs are not known to provide this partition and broadcast (Li et al., 2018) optimization by default.
3.2. Small-Large Outer-Joins
@@
For a left-outer-join, the Duplication and Efficient Redistribution (DER) algorithm (Xu and Kostamaa, 2010) broadcasts the small relation, $\mathcal{S}$, to all $n$ executors. On each executor, $e_{i}$, $\mathcal{S}$ is inner joined with $\mathcal{R}_{i}$. The joined records are output, and the ids of the unjoined records of $\mathcal{S}$ are distributed to the $n$ executors based on their hash. If an executor receives $n$ copies of a record id, then this $\mathcal{S}$ id is unjoinable with $\mathcal{R}$ (since it was unjoined on the $n$ executors). The unjoinable ids are inner Hash-Joined with $\mathcal{S}$ and $null$-padded. The left-outer-join result is the union of the two inner join results.
The Duplication and Direct Redistribution (DDR) algorithm (Cheng et al., 2017) is similar to DER, but for the records failing the first inner join, each executor hashes the entire unjoined record, instead of only its id, over the network. Because entire records, instead of their ids, are hash-distributed, (a) the first join can be executed as an out-of-the-box left-outer-join, and (b) the second inner join is not needed. Extending DER and DDR to full-outer-joins is achieved by changing the first join to output the unjoined records on $\mathcal{R}_{i}$.
4. The Tree-Join Algorithm
@@
4.1. The Basic Tree-Join Algorithm
@@
4:The second set of lists are the chunks of $\mathcal{L}_{2}$.
@@
Figure 3. A schematic example for joining a specific key, $k_{1}$, using the basic Tree-Join algorithm. (a) The $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The $k_{1}$ joined lists in the initial joined index formed using groupByKey. Each of the joined lists are chunked into two sub-lists (to simplify the figure). (c) Each sub-list from $\mathcal{R}$ is produced with all the sub-lists from $\mathcal{S}$. These pairs of sublists can be processed by different executors in the next iteration. (b) and (c) together show the top two levels of the tree formed by the basic Tree-Join algorithm.
This tree of the execution iterations is illustrated in Fig. 3 for an example key, $k_{1}$. Fig. 3(a) shows the $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. Fig. 3(b) shows the basic Tree-Join algorithm forming the joined lists of $k_{1}$ in the initial joined index, $joined\_index_{0}$, which is the top level of the tree. The joined lists are chunked into sub-lists. To simplify Fig. 3, each list is chunked into two sub-lists. That is, $\mathcal{L}_{\mathcal{R}1}$ is chunked into $\mathcal{L}_{\mathcal{R}11}$ and $\mathcal{L}_{\mathcal{R}12}$, and $\mathcal{L}_{\mathcal{S}1}$ is chunked into $\mathcal{L}_{\mathcal{S}11}$ and $\mathcal{L}_{\mathcal{S}12}$. Fig. 3(c) shows the basic Tree-Join algorithm forming the second joined index, $joined\_index_{1}$, which is the next level of the tree. This is done by outputting each sub-list from $\mathcal{R}$, $\mathcal{L}_{\mathcal{R}11}$ and $\mathcal{L}_{\mathcal{R}12}$, with all the sub-lists from $\mathcal{S}$, $\mathcal{L}_{\mathcal{S}11}$ and $\mathcal{L}_{\mathcal{S}12}$. Notice these joined sub-lists have a repeated key in $joined\_index_{1}$ Hence, these keys can be distributed on different data partitions and hence processed by different executors for load balancing. To simplify Fig. 3, only the first two levels of the execution iteration tree are shown in Fig. 3(b) and Fig. 3(c), respectively.
Figure 4. An example of inner-joining two relations using treeJoinBasic. (a) The input relations. (b) The initial index built from the input relations in (a). The initial index is used to produce (c) the first partial results, and (d) the first joined index. The first joined index is used to produce (e) the second partial results, and (f) the second joined index. The second joined index is used to produce (g) the third partial results. The union of the partial results in (c), (e), and (g) constitutes the inner-join results.
@@
4.2. Chunking Lists of Hot Keys
As explained in § 4.1, The treeJoinBasic algorithm executes multiple iterations. During each iteration, for each hot key, its joined lists are chunked by the splitter, the executor assigned the key and its joined lists, into sub-lists to be processed by the executors of the subsequent iteration. The subsequent executors can (a) output the pairs of records in the joined sub-lists, or (b) act as further splitters of the sub-lists if they are still too long. The chunking of joined sub-lists continues until they are short enough.
Computing the optimal number of sub-lists for a list of records is crucial for load balancing between the executors and across iterations. Chunking a pair of joined lists into numerous pairs of sub-lists adds overhead to the splitter. Conversely, chunking the pair of joined lists into a few pairs of sub-lists adds overhead to the executors that handle these pairs in the subsequent iteration.